If two charges -8q and+2q are separated by a distance equal to L, then where the third charge Q should be placed such that it is in equilibrium with the two charges? Plz reply.

Grade:11

1 Answers

Vikas TU
7 years ago
Force b/w -8q and Q would be:
F1 = k(8q)Q/x^2...............(1)
Force b/w +2q and Q would be:
F2 = k(2q)Q/(L +x)^2..............(2)
F1 = F2
8/x^2 = 2/(L+x)^2
4(L+x)^2 = x^2
(2L + 2x – x) (2L + 2x +x) = 0
x = -2L
and
x = -2L/3